### v1

# Mind the Gap. Doubling Constant Parametrization of Weighted Problems: TSP, Max-Cut, and More

Abstract
Despite much research, hard weighted problems still resist super-polynomial improvements over their textbook solution. On the other hand, the unweighted versions of these problems have recently witnessed the sought-after speedups. Currently, the only way to repurpose the algorithm of the unweighted version for the weighted version is to employ a polynomial embedding of the input weights. This, however, introduces a pseudo-polynomial factor into the running time, which becomes impractical for arbitrarily weighted instances.
In this paper, we introduce a new way to repurpose the algorithm of the unweighted problem. Specifically, we show that the time complexity of several well-known NP-hard problems operating over the $(\min,+)$ and $(\max,+)$ semirings, such as TSP, Weighted Max-Cut, and Edge-Weighted $k$-Clique, is proportional to that of their unweighted versions when the set of input weights has small doubling. We achieve this by a meta-algorithm that converts the input weights into polynomially bounded integers using the recent constructive Freiman’s theorem by Randolph and Węgrzycki [ESA 2024] before applying the polynomial embedding.
1 Introduction
The renaissance witnessed in exact algorithm design over the last two decades has one of its roots in the quest to answer the following fundamental question:
Are textbook solutions of NP-hard problems the best we can hope for?
Indeed, there are many problems that resist improvements over the standard algorithm. Yet, we have arguments to be optimistic: The amount of work in this direction has led to impressive speedups for well-studied problems such as Max-Cut [maxcut-williams], Hamiltonicity [hamiltonicity, below_all_subsets1], Sched [below_all_subsets2], Independent Set [below_all_subsets3], Dominating Set [below_all_subsets3, below_all_subsets4], and Min $k$-Cut [min_k_cut]. To some extent, these speedups have been achieved mainly for what we will refer to as unweighted problems. The weighted versions of many of these problems are still stuck with the textbook solutions that accept only modest, mostly polynomial-time, improvements.
A concrete example is that of the Traveling Salesman Problem (TSP), which asks to find the shortest tour through $n$ cities. The textbook algorithm in $O(2^{n}n^{2})$-time remains, up to some modest polynomial improvements [chan_shaving_logs], the best up to date for the general setting [bellman_dp, held_karp_dp, nederlof_tsp]. In contrast, its unweighted counterpart, the Hamiltonicity problem, has been (non-trivially) sped up to $O^{*}(1.66^{n})$-time by Björklund [hamiltonicity]. [^1] The $O^{*}$-notation hides polynomial factors in the input size.
Interestingly, virtually every new result on the unweighted version of a weighted problem has its own theorem stating that the original problem can be solved in the time of its unweighted counterpart, yet with an extra multiplicative pseudo-polynomial dependence on the largest input weight; this is known as the “(polynomial) embedding technique”. Notably, this situation also concerns TSP: Björklund proved that Hamiltonicity can be leveraged to solve TSP by paying the price of a near-linear time-factor depending on the largest edge weight $W$ [hamiltonicity, Thm. 3]. This becomes impractical for arbitrarily large $W$. The same holds true for the weighted Max-Cut problem: its unweighted counterpart can be solved in $O^{*}(2^{\omega n/3})$-time [maxcut-williams, koivisto-correct], where $\omega<2.371339$ is the matrix multiplication exponent [curr-omega], while its weighted version has to “accept” the additional pseudo-polynomial factor. On the other hand, if the weights are polynomially bounded, then the problematic factor becomes polynomial, and the running time of the weighted version matches, in the $O^{*}$-sense, that of the unweighted version.
Research Question.
Given the ubiquity of the embedding technique, it is natural to ask whether the gap between the running times of the weighted and unweighted versions of a problem is overcome only in the regime of polynomially bounded weights. A negative answer would be consistent with the widespread adoption of the embedding approach, with its inherent pseudo-polynomial dependence [maxcut-williams, koivisto-correct, fsc, set_partitioning, hamiltonicity, saving_space, Cygan_book]. In contrast, a positive answer would re-open the door to a more systematic re-examination of the weighted vs. unweighted gap, in the spirit of the “Losing Weights by Gaining Edges” technique for node-weighted problems [losing-ws] or in terms of approximation guarantees [apx-ws].
1.1 Our Results
Somewhat intriguingly, we show in this work that the polynomially bounded weights setting is not the only one that closes the gap between the weighted and unweighted versions of a problem. The meta-algorithm we introduce applies to a wide spectrum of NP-hard combinatorial problems that operate over the $(\min,+)$ and $(\max,+)$ semirings. Namely, it takes as input a weighted problem instance which guarantees that its weight set has small doubling, and solves it in time proportional to that of its unweighted counterpart. The set of input weights $A$ is said to have small doubling if $|A+A|\leq\mathcal{C}|A|$ for some constant $\mathcal{C}>1$, referred to as the doubling constant.
To maintain notation consistency with prior work on unweighted problems [esa-paper], we prefix problem names with $\mathcal{C}$-” if the set of input weights has doubling constant $\mathcal{C}$. We also write $O_{\mathcal{C}}$ to note that we suppressed factors that depend only on $\mathcal{C}$. Under this notation, all of our results conform to the following template: {restatable}[]corollaryDCTSP If Hamiltonicity can be solved in time $T(n)$, then $\mathcal{C}$-TSP can be solved in time $O_{\mathcal{C}}^{*}(T(n))$. The applicability of our work goes beyond TSP. Indeed, we can consider any weighted NP-hard problem that operates over the $(\min,+)$ and $(\max,+)$ semirings, such as the Weighted Max-Cut, Edge-Weighted $k$-Clique, or Minimum Steiner Tree problems. [^2] We kindly refer the reader to the relevant section for the problem definitions. The only requirement is that the problem at hand satisfies a property $\phi$: its objective value is an additive combination of input weights, and it admits an (algebraic) algorithm as seen earlier; see Prop. 1 for a formal treatment.
To avoid redundancy across different problems, we present our results under the umbrella of a meta-algorithm. Namely, given a weighted problem $\mathbf{P}_{w}$ that satisfies the above property $\phi$, its small-doubling counterpart can be solved in time matching that of the unweighted version:
{restatable}
[Meta-algorithm]theoremMETATHM If problem $\mathbf{P}_{w}$ satisfies property $\phi$ and the unweighted version can be solved in time $O(T(n))$, then $\mathcal{C}$-$\mathbf{P}_{w}$ can be solved in time $O_{\mathcal{C}}^{*}(T(n))$.
The following corollaries follow:
{restatable}
[]corollaryDCMAXCUT If unweighted Max-Cut can be solved in time $T(n,m)$, then $\mathcal{C}$-Weighted Max-Cut can be solved in time $O_{\mathcal{C}}^{*}(T(n,m))$. {restatable}[]corollaryDCEWCLIQUE If $k$-Clique can be solved in time $T(n,m,k)$, then $\mathcal{C}$-Edge-Weighted $k$-Clique can be solved in time $O_{\mathcal{C}}^{*}(T(n,m,k))$. {restatable}[]corollaryDCSTEINER If the unweighted version of Minimum Steiner Tree can be solved in time $T(n,m,k)$, then $\mathcal{C}$-Minimum Steiner Tree can be solved in time $O_{\mathcal{C}}^{*}(T(n,m,k))$.
We outline in the following how to obtain the results. Throughout the paper, we refer to the algebraic algorithm that exhibits the pseudo-polynomial dependence on $W$ as the bounded-input algorithm.
1.2 Overview of Approach
Let us provide some intuition of our approach before we continue with the general framework.
Intuition.
Consider TSP as a running example. Recall that the bounded-input algorithm works fine as long as $W$ is small and the weights are in $[0,W]$. Intuitively, we can “hack” the algorithm to also work with weights that are a subset of a simple arithmetic progression $\{V,V+r,\ldots,V+W\cdot r\}$, particularly for large values of $V$ and $r$, namely: Map the weights to $[0,W]$ by subtracting $V$ and dividing by $r$, compute the optimal tour as before on $[0,W]$, and, finally, recover the original objective by scaling by $r$ and adding $nV$. Thus, as long as the arithmetic progression is small, the algorithm still works. Our meta-algorithm lifts this idea to weights lying in a small generalized arithmetic progression—a structure that captures small-doubling sets—as follows.
Meta-Algorithm.
Let $A$ denote the set of input weights of the problem at hand, such as the edge weights in TSP. The first observation is that the objective value of the problem is contained in a $q$-fold sumset of $A$, i.e., $A+\ldots+A$ with $A$ occurring $q$ times, where $q$ is problem-specific; for TSP we have $q=n$. The second observation is that if $A$ has small doubling, then the cardinality of this very $q$-fold sumset is polynomial w.r.t. $O_{\mathcal{C}}$. [^3] Noteworthy, Plünnecke’s inequality is too loose in this setting, leading to an exponential bound. This suggests that, in principle, we could run the bounded-input algorithm. However, the values may still be numerically as large as $\Theta(q\cdot\max A)$. The missing piece is a decomposition of the weights into constant-size coefficient-tuples coming from the generalized arithmetic progression $A$ is a subset of (cf. Freiman’s theorem [freiman1]).
To be able to apply this to any bounded-input algorithm out of the box, we must ensure that the coefficient representation preserves the total order of the original weights. We ensure this by designing an order-preserving monomorphism from the coefficient-tuple space into $\{0,\ldots,|A^{\prime}|-1\}$, where $qA\subseteq A^{\prime}$ and $|A^{\prime}|$ remains polynomial w.r.t. $O_{\mathcal{C}}$.
1.3 Related Work
We survey applications of additive combinatorics to algorithm design and the literature on the unweighted-weighted dichotomy, focusing on the polynomial embedding technique.
Additive Combinatorics in Algorithm Design.
Our work continues the effort to consider important results from additive combinatorics [tao_vu] as a new toolbox for algorithm design, a connection highlighted in works such as Ref. [trevisan_ac, ac_cs_view3, ac_cs_view1, ac_cs_view2]. One of the first seminal results is that of Chan and Lewenstein [chan-lewenstein], who used the (constructive) Balog-Szemerédi-Gowers (BSG) theorem to speed up the monotone variant of the min-plus convolution on small integers. Recently, this has been further used and extended in #APSP [fredman_meets_domi], 3SUM [3sum_sidon], listing 4-cycles [abf], and approximate distance oracles [abf]. We observe that the constructive version of a fundamental theorem in additive combinatorics bears fruit in algorithm design. To this end, Randolph and Węgrzycki [esa-paper] made Freiman’s theorem [freiman1], another fundamental result in additive combinatorics, constructive. Their main applications were to SubsetSum and ILP-Feasibility, which have an inherent additive structure. Indeed, we extend their results to the area of weighted problems.
Embedding Technique.
Repurposing the unweighted algorithm for the weighted version of a problem is a common theme in work that improves over the textbook solution, as in seminal results such as Ref. [maxcut-williams, fsc, set_partitioning, hamiltonicity]; notably, in all these examples, the weighted problem operates over the $(\min,+)$ and $(\max,+)$ semirings. The key idea is to represent the input weights of the problem instance as monomials so that the bounded-input algorithm can work in the $(+,\times)$ ring. The drawback is that this introduces a pseudo-polynomial dependence in the running time, since operations now incur an additional $\widetilde{O}(W)$ factor, [^4] The $\widetilde{O}-$notation hides polylogarithmic factors in the input size. where $W$ is the largest input weight. For instance, undirected Hamiltonicity runs in $O^{*}(1.66^{n})$-time, hence TSP can be solved in $O^{*}(1.66^{n}W)$-time. This, however, becomes impractical for arbitrarily large $W$. Still, if $W$ is polynomially bounded, then up to polynomial factors, the weighted problem becomes equivalent to the unweighted one. For example, if $W=\text{poly}(n)$, then TSP can be solved in time $O^{*}(1.66^{n})$, matching its unweighted counterpart.
Dodging Pseudo-Polynomiality.
The pseudo-polynomial factor is often unwanted. One way to shave it is to enter the realm of approximation. In fact, there is a simple exercise to turn any bounded-input algorithm for a problem into an exponential-time $(1+\varepsilon)$-approximation algorithm. For the Max-Cut problem, Williams mentions this idea in passing [maxcut-williams], inspired by the literature on the all-pairs shortest paths problem [zwick_apsp]; see, e.g., Ref. [approx-min-sum-conv] for more such examples. It is interesting to ask whether the doubling constant could play a role in approximation algorithms.
1.4 Roadmap
We structure the paper as follows. We start with preliminaries on the embedding technique (Sec. 2.1) and the necessary notation in additive combinatorics (Sec. 2.2). Then, we introduce the meta-algorithm (Sec. 3), including the order-preserving monomorphism (Sec. 3.1). Subsequently, we instantiate the meta-algorithm on several weighted problems (Sec. 4). We conclude with an outlook on polynomial-time weighted problems (Sec. 5).
2 Preliminaries
Notation.
For $a<b\in\mathbb{Z}$, define $a,b]\triangleq\{a,\ldots,b\}$. As shorthand, we define $n]\triangleq[1,n]$. Moreover, we define $f(X)$ as $\{f(x)\mid x\in X\}$ for a function $f$ and set $X$.
Time Complexities.
As we will only deal with exponential-time algorithms, we use the $O^{*}$-notation to hide polynomial factors in the input size. For example, $O^{*}(2^{n})=O(2^{n}\text{poly}(n))$. In particular, this notation will never hide pseudo-polynomial factors depending on the input weights; specifically, we mean the largest input weight, which will be denoted by $W$. Furthermore, we use the $O_{\mathcal{C}}$-notation, already established in Ref. [esa-paper], to denote that we suppressed factors depending on $\mathcal{C}$, the doubling constant on which we parameterize the time-complexities. For instance, $O_{\mathcal{C}}(2^{n})=f(\mathcal{C})\cdot O(2^{n})$ for a computable function $f$. To put into context, $O_{\mathcal{C}}^{*}(2^{n})$ will thus mean $f(\mathcal{C})\cdot O^{*}(2^{n})$.
Graphs.
Unless otherwise specified, all problems we consider in this work operate on an undirected graph $G=(V,E)$, most of the time endowed with a weight function $w$ defined on the edge set $E(G)$ that attributes each edge $e$ a positive value. The length of a path $\pi=(v_{1},\ldots,v_{k})$ is k1k-1, i.e., we count the number of edges, while the weight of the path is the sum of all edge weights, i.e., $w(\pi)\vcentcolon=\sum_{i=1}^{k-1}w(v_{i},v_{i+1})$.
2.1 Embedding Technique
We present the embedding technique, which is used by several works to leverage the running time of the unweighted problem for the corresponding weighted counterpart, at the cost of a pseudo-polynomial dependence on the largest input weight, $W$. Specifically, we consider NP-hard problems that admit formulations over the $(\min,+)$ and $(\max,+)$ semirings: combining substructures corresponds to addition, while selecting among candidate values corresponds to taking a minimum (for minimization) / maximum (for maximization).
Overview.
The key idea is to encode weights as exponents in a polynomial ring, since (i) the “++” operator of our semirings translates to multiplication between monomials ($x^{a+b}=x^{a}\cdot x^{b}$), and (ii) aggregating candidate values corresponds to polynomial addition, which retains information about all previously attainable values. [^5] If the reader will, this operation “simulates” an additive inverse. Hence, the output of the bounded-input algorithm is a solution polynomial:
Definition 2.1 (Solution polynomial).
Let $I_{w}$ be an instance of a weighted combinatorial problem with a finite set $\mathcal{F}(I_{w})$ of feasible solutions. For each $\sigma\in\mathcal{F}(I_{w})$, let $v(\sigma)\in\mathbb{Z}_{\geq 0}$ denote its objective value. The solution polynomial of $I_{w}$ is the ordinary generating function
$F_{I_{w}}(x)=\sum_{\sigma\in\mathcal{F}(I_{w})}x^{v(\sigma)}=\sum_{s=0}^{V_{\max}}m_{I_{w}}(s)\,x^{s},$
where $m_{I_{w}}(s)\coloneqq\bigl|\{\sigma\in\mathcal{F}(I_{w})\mid v(\sigma)=s\}\bigr|$ and $V_{\max}\coloneqq\max_{\sigma\in\mathcal{F}(I_{w})}v(\sigma)$.
The solution polynomial abstracts away whether the problem is a minimization or maximization problem: for minimization we select the monomial with the smallest exponent, whereas for maximization we select the monomial with the largest exponent. To this end, we refer to an algorithm as algebraic, if it returns a solution polynomial for an instance of a problem. [^6] An alternative attribute could be non-combinatorial. For a treatment on the definition of a combinatorial algorithm, see Ref. [fast-bmm].
Due to the additive structure of the input weights we will be exploiting, starting in Sec. 3, we will have to work with coordinate tuples instead of the standard setting of integer weights. To overcome this, we will propose a pairing function that transforms tuples into equivalent integer weights; see Sec. 3.1 for the technical details.
2.2 Additive Combinatorics
As the trend of importing results from additive combinatorics into algorithm design is itself rather young, we provide the necessary notation so that our algorithms can be easily followed.
Basic Notation.
Let $A$ and $B$ be two sets. Define their sumset (also known as the Minkowski sum) $A+B$ in a commutative group as $\{a+b\mid a\in A,b\in B\}$. When referring to $A+A$, we directly note $2A$. Hence, the iterated sumset $h$-fold hAhA is defined recursively by hA=(h1)A+AhA=(h-1)A+A.
Doubling Constant.
An important definition is that of the doubling constant of a set $A$, which quantifies how large the sumset becomes w.r.t. to the original set:
$\mathcal{C}(A)=\frac{|A+A|}{|A|}.$ (1)
Whenever $\mathcal{C}(A)$ does not depend on the size of $A$, we directly write $\mathcal{C}$. We say that these sets have small doubling. Our work shows that whenever the set of input weights of NP-hard problems has small doubling, that instance can be solved faster.
This comes as a surprise given the simple definition of the doubling constant in Eq. (1). To gain an intuitive understanding, consider the case when $A$ is a (simple) arithmetic progression, say $\{2,4,6,8\}$. Its 2-fold sumset, $A+A$, is $\{4,6,8,10,12,14,16\}$. Taking a more general example will lead us to conclude that the sumset of a (simple) arithmetic progression does not explode that much. Indeed, |A+A|=2|A|1|A+A|=2|A|-1, whenever $A$ is a (simple) arithmetic progression. On the contrary, the set $\{3,5,9,17\}$ results in a sumset that attains the maximum size for a 4-size set, namely 10 elements. In other words, there are no distinct pairs in $A\times A$ that have the same sum.
GAPs.
Apart from the well-known simple arithmetic progression, additive combinatorics results target generalized arithmetic progressions (GAP). Formally, a GAP $G$ is defined as
$G=\{x_{1}\ell_{1}+\ldots+x_{d}\ell_{d}\mid\forall i\in[d],\ell_{i}\in[0,L_{i}]\},$
where $d$ is the dimension of the GAP, $\{x_{1},\ldots,x_{d}\}$ are the generators, and $\{L_{1},\ldots,L_{d}\}$ are the dimension bounds; in particular, a 1-dimensional GAP is a simple arithmetic progression.
In our previous examples, we saw that if $A$ is an arithmetic progression, then it has a small doubling constant; the interesting problems in additive combinatorics are indeed the inverse problems: What does a set of small doubling look like? This is where Freiman’s theorem comes into play.
Freiman’s Theorem.
Given a set $A$ of small doubling, Freiman’s theorem [freiman1] tells that, indeed, $A$ is included in a GAP $G$, whose dimension is independent on $|A|$. In this work, we will need the constructive Freiman’s theorem, recently designed by Randolph and Węgrzycki [esa-paper]:
Theorem 2.2 (Constructive Freiman’s Theorem [esa-paper]).
Let $A$ be a set of $n$ integers with $|A+A|\leq\mathcal{C}|A|$. Then, there exists an $\tilde{O}_{\mathcal{C}}(n)$-time algorithm that with probability $1-n^{-\gamma}$, for an arbitrarily large constant $\gamma>0$, returns a generalized arithmetic progression
$G=\{x_{1}\ell_{1}+x_{2}\ell_{2}+\dots+x_{d(\mathcal{C})}\ell_{d(\mathcal{C})}\mid\forall i,\ell_{i}\in[L_{i}]\}\supseteq A$
with dimension $d(\mathcal{C})$ and volume $v(\mathcal{C})|A|$, where $d$ and $v$ are computable functions that depend only on the doubling constant $\mathcal{C}$. Specifically, the theorem provides the values $x_{1},x_{2},\dots,x_{d(\mathcal{C})}$ and $L_{1},L_{2},\dots L_{d(\mathcal{C})}$.
The above theorem gives us a GAP of dimension $d(\mathcal{C})=2^{\mathcal{C}^{O(1)}}$ and volume $v(\mathcal{C})=2^{2^{\mathcal{C}^{O(1)}}}n$, as in the original Freiman’s theorem [freiman1]. Noteworthy, later works do optimize $d$ and $v$ [freiman-better-1, freiman-better-2, freiman-better-3, freiman-better-4], yet these improvements are not captured by the constructive version.
This completes the necessary preliminaries on additive combinatorics. We are now ready to show that, in the $O_{\mathcal{C}}$-sense, NP-hard weighted problems reduce to their unweighted versions whenever their input weights form a set of small doubling.
3 Meta-Algorithm
Before outlining the meta-algorithm, we first define the property that the considered weighted problems have to satisfy to be amenable to the recent speedups achieved for their unweighted counterparts. As motivated in the introduction, the key requirement is that the objective value has to be an additive combination of input weights. This additive structure ensures that objective values lie within a folded sumset of the input weights.
Property 1 (Property $\phi$).
Let $\mathbf{P}_{w}$ be a weighted problem and $\mathbf{P}$ its unweighted counterpart. Let $I$ be an instance of $\mathbf{P}_{w}$ of size $n$, $w(I)$ its set of weights, and $W=\max w(I)$. Then, we say that $\mathbf{P}_{w}$ has property $\phi$ if the following hold:
  1. 1.
    Any feasible solution to $I$ is the total weight of a set $S\subseteq w(I)$,
  2. 2.
    There is an algebraic algorithm $\mathcal{A}$ that solves $\mathbf{P}$ in $O(T(n))$-time and $\mathbf{P}_{w}$ in $O^{*}(T(n)\cdot W)$-time, and any intermediate solution to $I$ produced by $\mathcal{A}$ is the total weight of a polynomial-size multi-set with support in $w(I)$.
Notably, the additional requirement in Condition 2 is inherent to any bounded-input algorithm whose running time has the near-linear dependence on the largest input weight $W$. Next, we introduce the pairing function that maps GAP coordinates to integers, allowing the bounded-input algorithm to operate directly on integers rather than raw coordinate tuples.
3.1 Pairing Function
To be able to actually run the bounded-input algorithm $\mathcal{A}$, we need to transform GAP coordinates into integers. In particular, the meta-algorithm will first enlarge the dimension bounds before encoding them as integers, so that the coordinates set is well-defined, namely: Given any two GAP coordinate tuples $\alpha=\langle\alpha_{1},\ldots,\alpha_{d}\rangle$ and $\beta=\langle\beta_{1},\ldots,\beta_{d}\rangle$ of dimension $d$, it holds that $\alpha_{i}+\beta_{i}\leq L_{i}$, $\forall i\in[d]$.
In addition, as we operate over the $(\min,+)$ and $(\max,+)$ semirings, the pairing function $\kappa$ has to preserve the “additivity” of the coordinates, namely: [^7] Note that we skip redundant information from $\kappa$, such as the dimension and the dimension bounds.
$\kappa(\alpha\oplus\beta)=\kappa(\alpha)+\kappa(\beta),$ (2)
where $\oplus$” is the entrywise addition operator, i.e., $\langle\alpha\oplus\beta\rangle_{i}\vcentcolon=\alpha_{i}+\beta_{i},\forall i\in[d]$. Since $\kappa$ must be injective, so that we can uniquely get back the GAP coordinates, $\kappa$ has to be a monomorphism between a well-defined set of GAP coordinates and the encoded integers.
Our Pairing Function.
We propose the following pairing function:
$\kappa\left(d,\langle L_{i}\rangle_{i\in[d]},\langle l_{i}\rangle_{i\in[d]}\right)=\begin{cases}l_{1},&\text{if }d=1,\\ l_{d}+(L_{d}+1)\cdot\kappa\left(d-1,\langle L_{i}\rangle_{i\in[d-1]},\langle l_{i}\rangle_{i\in[d-1]}\right)\!,&\text{if }d>1.\end{cases}$
It receives the dimension of the GAP, $d$, its dimension bounds $\langle L_{i}\rangle_{i\in[d]}$, and the GAP coordinates $\langle l_{i}\rangle_{i\in[d]}$ of an element, and returns the encoded value. If there is only one dimension, then the corresponding GAP coordinate is immediately returned. Otherwise, it builds the encoded value recursively, by multiplying the result for the remaining d1d-1 dimensions by $L_{d}+1$ and finally adding $l_{d}$. We show that in our setting it holds that $\kappa$ is a monomorphism:
{restatable}
[]lemmaPAIRING Let $\langle L_{i}\rangle_{i\in[d]}\in\mathbb{N}^{d}$ be the bounds of a GAP of dimension $d$. Given a well-defined set $D\subseteq\prod_{i=1}^{d}[0,L_{i}]$ of $d$-size coordinate tuples, the pairing function $\kappa$ is a monomorphism between $(D,\oplus)$ and $([0,\prod_{i=1}^{d}(L_{i}+1)-1],+)$. We defer the proof of Lemma 3.1 to Appendix A. As a by-product of our injectivity proof, we show that $\kappa$ satisfies a certain monotonicity property:
$\alpha\prec\beta\implies\kappa(\alpha)<\kappa(\beta),$ (3)
where $\alpha\prec\beta$ stands for the (syntactic) lexicographical order of the coordinate tuples, e.g., $\langle 1,2\rangle\prec\langle 1,3\rangle,\langle 1,2\rangle\prec\langle 2,1\rangle$. However, even then, there is still an issue: monomorphism alone does not suffice. We discuss why this is the case in the following.
Why Monomorphism Alone Is Not Sufficient.
[…]
Restoring the Order.
To address this, we augment the call to the bounded-input algorithm with a (fixed) permutation $\pi$ that restores the order of the original values. Whenever the algorithm has to find the minimum / maximum monomial in the solution polynomial (Def. 2.1)—the final last step in the bounded-input algorithms we consider—, it uses $\pi$ to recover the true rank of the current encoding $e$.
Building $\pi$.
The permutation can be built in polynomial-time w.r.t. $O_{\mathcal{C}}$, as follows: Once the dimension bounds of the GAP have been enlarged (see line 3 in the upcoming Alg. 3.2), we can generate the values $\sum_{i=1}^{d}x_{i}\cdot l_{i}$ for $l_{i}\in[0,\lambda L_{i}],\forall i\in[d]$, where $\lambda$ is problem-specific (see the upcoming Sec. 3.2). By construction, these are all possible values that can be taken during a bounded-input algorithm’s run. To actually “freeze” the permutation $\pi$, we have to sort the generated values, which takes time $\widetilde{O}(\lambda^{d}\prod_{i=1}^{d}(L_{i}+1))$. We call this step BuildPermutation in the meta-algorithm.
Decoding.
The decoding process follows a similar line and is outlined in the following:
$\kappa^{-1}\left(d,\langle L_{i}\rangle_{i\in[d]},e\right)=\begin{cases}\langle e\rangle,&\text{if }d=1,\\ \kappa^{-1}\left(d-1,\langle L_{i}\rangle_{i\in[d-1]},\left\lfloor\frac{e}{L_{d}+1}\right\rfloor\right)\circ\langle e\bmod{(L_{d}+1)}\rangle,&\text{if }d>1.\end{cases}$
Namely, if there is only one dimension, the current encoding is returned, otherwise we append the result of $e\bmod{(L_{d}+1)}$ to the tuple resulting from the remaining d1d-1 dimensions. The recursive call is done with an $e$ from which the contribution of dimension $d$ has been removed.
3.2 Exploiting the Bounded-Input Algorithm
Let $\mathbf{P}_{w}$ be a weighted problem satisfying property $\phi$, and let $I$ be an instance of $\mathbf{P}_{w}$ of size $n$ whose set of weights has small doubling. Let $\mathcal{A}$ be the bounded-input algorithm from Condition 2. Furthermore, let $\lambda=\text{poly}(|w(I)|)$ be an upper-bound on the cardinality of the multi-sets from the same condition; as we will see, for all bounded-algorithms we consider in this paper it holds that $\lambda=O(|w(I)|)$. We can now state our meta-algorithm, outlined in Alg. 3.2.
{restatable}
[]algorithmMETA_ALGO $\mathcal{M}(I,\mathcal{A})$
[…]
7:return $\sum_{i=1}^{d(\mathcal{C})}x_{i}l_{i}^{\prime}\>(=\vcentcolon c^{*})$
[…]
Proof.
We prove that the bounded-input algorithm actually receives as input polynomially bounded weights w.r.t. $O_{\mathcal{C}}$ via the new weight function $w^{\prime}$ (line 5). This will guarantee that the line runs in time $O_{\mathcal{C}}^{*}(T(n))$. Note that this does not hide pseudo-polynomial factors in the largest input weight.
To this end, let $G^{\prime}$ be the GAP which $\lambda w(I)$, a superset of any intermediate solution—recall Condition 2, is a subset of. Moreover, the initial set of input weights $w(I)$ is also a subset of a GAP $G$. Let us see the connection between $G^{\prime}$ and the original $G$. First, the parameters of $G$ are $\{x_{1},\ldots,x_{d(\mathcal{C})}\}$ and $\{L_{1},\ldots,L_{d(\mathcal{C})}\}$, i.e.,
$G=\{x_{1}l_{1}+\ldots+x_{d(\mathcal{C})}l_{d(\mathcal{C})}\mid\forall i,l_{i}\in[L_{i}]\}.$
Then, since $G^{\prime}$ has to capture the weights in $\lambda w(I)$, we need to modify the bounds of the original GAP. We thus obtain the following new GAP:
$G^{\prime}=\{x_{1}l_{1}+\ldots+x_{d(\mathcal{C})}l_{d(\mathcal{C})}\mid\forall i,l_{i}\in[\lambda L_{i}]\}.$
Recall now that the coordinates are mapped to the integer range $0,\prod_{i=1}^{d(\mathcal{C})}(\lambda L_{i}+1)-1]$, which is exactly the volume of $G^{\prime}$. Bounding this, we obtain:
$|G^{\prime}|=\prod_{i=1}^{d(\mathcal{C})}(\lambda L_{i}+1)$ $\leq\lambda^{d(\mathcal{C})}\prod_{i=1}^{d(\mathcal{C})}(L_{i}+1)$
$=\lambda^{d(\mathcal{C})}|G|$
$=\lambda^{d(\mathcal{C})}v(\mathcal{C})\,|w(I)|$
$=|w(I)|^{O_{\mathcal{C}}(1)}.$ (4)
Therefore, the bounded-input algorithm (line 5) is run on an instance with polynomially large weights w.r.t. $O_{\mathcal{C}}$, represented by the weight function $w^{\prime}$.
Let us analyze the total running time. We argue that all steps apart from line 5 take polynomial time w.r.t. $O_{\mathcal{C}}$ each. Namely, once the parameters of the GAP $G$ have been computed (line 1), we can compute the coordinates corresponding to the values of the weight function $w$ by a naive iteration of $G$ (line 2); this takes time $O(|w(I)|\cdot|G|)=O(|w(I)|\cdot v(\mathcal{C})|w(I)|)=O_{\mathcal{C}}(|w(I)|^{2})$. Then, building the permutation $\pi$ (line 4) takes time $\widetilde{O}(|G^{\prime}|)=\widetilde{O}(|w(I)|^{O_{\mathcal{C}}(1)})$; see Sec. 3.1 and the above Eq. (4). Finally, applying $\kappa$ on the GAP coordinates $w_{\text{coord}}$ takes $O_{\mathcal{C}}(|w(I)|)$-time (line 3), while lines 6 and 7 take $O_{\mathcal{C}}(1)$-time and $\widetilde{O}_{\mathcal{C}}(1)$-time, respectively.
Noteworthy, when not parameterizing on the doubling constant, the running time of the meta-algorithm reads $O^{*}(\lambda^{d(\mathcal{C})}v(\mathcal{C})T(n))$, which is $O_{\mathcal{C}}^{*}(T(n))$. ∎
4 Applications
In the following, we review several weighted NP-hard problems and show how to solve their small-doubling instances faster than the traditional algorithm by leveraging the previous meta-algorithm. This enables running times in the regime of the unweighted version of the problem for a large suite of problems, such as TSP, Weighted Max-Cut, Edge-Weighted $k$-Clique, and Minimum Steiner Tree. We stress that, compared to prior work [esa-paper], we consider weighted problems that operate over tropical semirings.
4.1 TSP Meets Small Doubling
We define the doubling constant parametrization of TSP, $\mathcal{C}$-TSP, as follows:
$\mathbf{\mathcal{C}}$-TSP
Input: Undirected graph $G=(V,E)$, a cost $w(e)\in\mathbb{Z}_{\geq 0}$ for each $e\in E=E(G)$, such that $|w(E)+w(E)|\leq\mathcal{C}|w(E)|$ Output: Tour $T\subseteq E$ that minimizes $\sum_{e\in T}w(e)$
Why the Embedding Technique Fails.
First, note that bounded-input algorithm does not help much for $\mathcal{C}$-TSP: The largest edge weight $W$ can still be exponentially large. Therefore, the resulting running time would be even worse than that of the traditional algorithm. We show an alternative solution via our meta-algorithm (Alg. 3.2).
Intuition.
Let $A\vcentcolon=w(E)$ denote the set of edge weights. Recall that the weight of any tour in TSP is computed by the total sum of the edge weights on that tour. To understand how this actually relates to the doubling constant, consider the case when we only consider paths of length 2. The possible weights of these paths are all included in $A+A$. Note that not all path weights are also valid: some weights might come from paths using the same edge twice. This, however, is not an issue. The main benefit we get is that the cardinality of $A+A$ is at most $\mathcal{C}|A|$. If we continue with longer paths, we will observe the same situation: Namely, the weights of the paths of length $k$ will be included in $kA$, the $k$-fold sumset of $A$. In particular, the shortest tour length will be an element of $nA$.
Lemma 4.1.
TSP satisfies property $\phi$.
Proof.
Condition 1 holds since a tour’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Björklund’s algorithm [hamiltonicity] for undirected Hamiltonicity, which runs in time $O^{*}(1.66^{n})$, and solves the weighted case in time $O^{*}(1.66^{n}\,W)$ [hamiltonicity, Thm. 3]. Due to the algorithm’s design, namely that it counts cycle covers of $n$ arcs, we have $\lambda\leq n$. ∎
[…]
4.2 Weighted Max-Cut Under Small Doubling
Another well-studied problem whose unweighted version has been sped up and fits the small doubling setting well is the Weighted Max-Cut problem. In a breakthrough result, Williams [maxcut-williams] improved the time complexity of the unweighted Max-Cut to $O^{*}(2^{\omega n/3})$. We show that we can leverage this result to the weighted setting when the edge weights form an integer set that has small doubling. We formally define the doubling constant parametrization of the problem in the following:
$\mathbf{\mathcal{C}}$-Weighted Max Cut
Input: Undirected graph $G=(V,E)$, a weight $w(e)\in\mathbb{Z}_{\geq 0}$ for each $e\in E=E(G)$ such that $|w(E)+w(E)|\leq\mathcal{C}|w(E)|$ Output: Subset of vertices $S$ so that $\sum_{e\in\delta(S)}w(e)$ is maximized
Lemma 4.2.
Weighted Max-Cut satisfies property $\phi$.
Proof.
Condition 1 holds since a cut’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Williams’s algebraic algorithm [maxcut-williams] (and later improved by Koivisto [koivisto-correct]) for the unweighted Max-Cut problem, which runs in time $O^{*}(2^{\omega n/3})$, and solves the weighted case in time $O^{*}(2^{\omega n/3}\,W)$ [koivisto-correct, Thm. 3]. ∎
Using the above lemma together with Thm. 1.1, we obtain:
[…]
4.3 Small Doubling and Edge-Weighted $k$-Cliques
We now turn to a problem whose complexity is a rather popular hypothesis in hardness proofs [clique-hyp-1, clique-hyp-2, clique-hyp-3]: the edge-weighted $k$-clique problem. Compared to the problems considered so far, the algebraic algorithm for the unweighted $k$-Clique has already been known from the 80’s [uw-k-clique]. We show that it can be leveraged in the doubling constant parametrization of the edge-weighted version. Note that, apart from a mention in passing in Lincoln et al. [clique-hyp-3] that Nešetřil-Poljak’s algorithm can also be used for polynomially bounded weights, we could not find any reference to a proof. For completeness, we show this in Lemma B (Appendix B).
The doubling constant parametrization of the problem is the following:
$\mathbf{\mathcal{C}}$-Edge-Weighted $k$-Clique
Input: Undirected graph $G=(V,E)$, integer $k$, and a weight $w(e)\in\mathbb{Z}_{\geq 0}$ for each $e\in E=E(G)$ such that $|w(E)+w(E)|\leq\mathcal{C}\,|w(E)|$ Output: A vertex set $S\subseteq V$ with $|S|=k$ that induces a clique and maximizes $\sum_{e\in E(G[S])}w(e)$
Lemma 4.3.
Edge-Weighted $k$-Clique satisfies property $\phi$.
Proof.
Condition 1 holds since a clique’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as Nešetřil-Poljak’s algebraic algorithm [uw-k-clique] for the unweighted version of the problem (or the improvement by Eisenbrand and Grandoni [faster-uw-k-clique] for some values of $k$ using rectangular matrix multiplication), which runs in time $O^{*}(n^{k\omega/3})$, and solves the weighted case in time $O^{*}(n^{k\omega/3}\,W)$ (Lemma B). Notably, due to its design, namely that each original input weight is doubled in the auxiliary graph, the bounded-input algorithm (Lemma B) has a $\lambda\leq 2|w(I)|$. ∎
In combination with Thm. 1.1, we obtain:
[…]
4.4 Steiner Meets Freiman
For reference, we also include the Minimum Steiner Tree, even though the general setting can be directly solved in $O((2+\epsilon)^{k}\,p(n))$ [faster_steiner1], where $p(n)$ is a polynomial function of $n$, the degree of which grows rapidly as $\epsilon$ approaches zero, Its degree has been later refined to $12\sqrt{\epsilon^{-1}\ln\epsilon^{-1}}$, having as a result running times such as $O(2.1^{k}\,n^{57.6})$ or $O(2.5^{k}\,n^{14.2})$ [fuchs2007dynamic]. Notably, Dreyfus and Wagner [dreyfus1971steiner] were the first to design a dynamic programming recursion running in time $O(3^{k}n+2^{k}n^{2}+nm)$, where $n=|V|$, $m=|E|$, and $k=|T|$.
The doubling constant parametrization of the problem is the following:
$\mathbf{\mathcal{C}}$-Minimum Steiner Tree
Input: Undirected graph $G=(V,E)$, a weight $w(e)\in\mathbb{Z}_{\geq 0}$ for each $e\in E=E(G)$, and a set of vertices $K\subseteq V=V(G)$ Output: Subgraph $H$ of $G$ that connects the vertices in $K$ and has the minimum total weight $\sum_{e\in E(H)}w(e)$ among all such subgraphs of $G$
Next, we show that the minimum Steiner tree is amenable for speedups in this regime:
Lemma 4.4.
Minimum Steiner Tree satisfies property $\phi$.
Proof.
Condition 1 holds since, as the edge weights are all non-negative, the optimal subgraph $H$ is a tree with its leaves in $K$, and a tree’s weight is the sum of the weights of its constituent edges. Next, we can take $\mathcal{A}$ as the fast subset convolution based algorithm due to Björklund, Husfeldt, Kaski, and Koivisto [fsc, Sec. 4.1.2], which runs in $O^{*}(2^{k}n^{2}W)$-time, or Lokshtanov–Nederlof’s polynomial-space algorithm [saving_space] in the same running time. ∎
With this, using Thm. 1.1, we obtain:
[…]
This concludes the suite of applications of our meta-algorithm. We conclude with an outlook on the applicability of our framework beyond NP-hard weighted problems.
5 Outlook: Beyond NP-hard Problems
As shown in this paper, the constructive Freiman’s theorem can be converted in a rather effective application for NP-hard weighted problems as well. A natural question arises: Is it possible to extend its applicability to weighted problems in P?
Min-Plus Convolution.
The answer seems to be a positive one, in particular for those polynomial problems for which the solution is an additive combination of the input weights. A simple application is that of the $(\min,+)$-convolution: Given sequences $(a[i])_{i=0}^{n-1}$ and $(b[i])_{i=0}^{n-1}$, the goal is to compute $(c[i])_{i=0}^{n-1}$, where $c[k]=\min_{i=0,\ldots,k}\{a[i]\,+\,b[k-i]\}$. The naive algorithm runs in $O(n^{2})$-time, and no $O(n^{2-\varepsilon})$-time algorithm is known for $\varepsilon>0$ [min-plus-1, min-plus-2]. However, the convolution in the $(+,\times)$-ring can be solved in $O(n\log n)$-time via FFT. This resembles the setting we considered throughout the paper. Indeed, Węgrzycki outlines in his PhD thesis a bounded-input algorithm running in time $\widetilde{O}(nW)$, where $W$ is the largest input value [karol_phd, Lemma 5.7.2]. Applying the same key idea as for the NP-hard problems we considered, we can obtain an $\widetilde{O}_{\mathcal{C}}(n)$-time algorithm for the min-plus self-convolution problem, i.e., $a=b$, and the input sequence, when regarded as a set, has small doubling. The same setting holds true for the min-sum subset convolution [fsc].
It is thus interesting to ask how the framework behaves in more weighted applications, such as (specialized) $(\min,+)$ matrix products [special-min-plus-prod] or even APSP, where—similar to the NP-hard case studied here—the objective value is an additive combination of input weights.
Acknowledgments.
The author thanks the anonymous reviewers of STACS’25 and STACS’26, whose detailed comments visibly improved the quality of the presentation.
Appendix A Pairing Function’s Monomorphism
[…]
Proof.
We first prove that $\kappa$ is a homomorphism and then prove that it is also injective.
Homomorphism. Let $\alpha=\langle\alpha_{1},\ldots,\alpha_{d}\rangle$ and $\beta=\langle\beta_{1},\ldots,\beta_{d}\rangle$ be two GAP-coordinates. Their entrywise addition $\alpha\oplus\beta$ is defined as
$\alpha\oplus\beta=\langle\alpha_{1}+\beta_{1},\ldots,\alpha_{d}+\beta_{d}\rangle.$
At this point, we use the fact that $D$ is well-defined, i.e., the dimension bounds are not exceeded: $\alpha_{i}+\beta_{i}\leq L_{i},\forall i\in[d]$. (This is guaranteed by the fact that we enlarge the dimension bounds before running the bounded-input algorithm; compare line 3 in Alg. 3.2). Formally, we have to prove that
$\kappa\left(d,\langle L_{i}\rangle_{i\in[d]},\alpha\oplus\beta\right)=\kappa\left(d,\langle L_{i}\rangle_{i\in[d]},\alpha\right)+\kappa\left(d,\langle L_{i}\rangle_{i\in[d]},\beta\right).$ (5)
In the sequel, we use the notation $\kappa_{d-1}(x)$ to denote that $\kappa$ is applied only on the first d1d-1 dimensions of a tuple of GAP coordinates $x$, i.e., $\kappa\left(d-1,\langle L_{i}\rangle_{i\in[d-1]},\langle x_{i}\rangle_{i\in[d-1]}\right)$; analogously, $\kappa_{d}$ will refer to the first $d$ dimensions. To prove Eq. (5), we use induction on the number of dimensions. The base case, $d=1$, follows by construction. Otherwise, assume Eq. (5) holds for d1d-1 dimensions (IH1), i.e.,
$\kappa_{d-1}(\alpha\oplus\beta)=\kappa_{d-1}(\alpha)+\kappa_{d-1}(\beta).$
Then,
$\kappa_{d}\left(\alpha\oplus\beta\right)$ $=\kappa_{d}\left(\alpha\right)+\kappa_{d}\left(\beta\right)$
$\overset{\kappa}{\iff}$ $(\alpha\oplus\beta)_{d}+(L_{d}+1)\kappa_{d-1}\left(\alpha\oplus\beta\right)$ $=\alpha_{d}+\beta_{d}+(L_{d}+1)\left(\kappa_{d-1}(\alpha)+\kappa_{d-1}(\beta)\right)$
$\overset{\text{IH1}}{\iff}$ $(\alpha\oplus\beta)_{d}$ $=\alpha_{d}+\beta_{d},$
which is exactly the definition of $\alpha\oplus\beta$
Injectivity. We next prove the injectivity of $\kappa$, by showing that $\kappa$ actually satisfies a certain monotonicity property, outlined in the following:
$\alpha\prec\beta\implies\kappa_{d}(\alpha)<\kappa_{d}(\beta),$ (6)
where $\alpha\prec\beta$” means that $\alpha$ is strictly lexicographically smaller than $\beta$. We prove this property by induction on the number of dimensions. The base case, $d=1$ follows by construction. Now, assume that property Eq. (6) holds for the first d1d-1 dimensions (IH2), i.e.,
$\langle\alpha_{i}\rangle_{i\in[d-1]}\prec\langle\beta_{i}\rangle_{i\in[d-1]}\implies\kappa_{d-1}(\alpha)<\kappa_{d-1}(\beta).$
To prove the original Eq. (6), we differentiate between two cases when $\alpha\prec\beta$ holds true:
Case
$\langle\alpha_{i}\rangle_{i\in[d-1]}=\langle\beta_{i}\rangle_{i\in[d-1]}$. This means that $\kappa_{d-1}(\alpha)=\kappa_{d-1}(\beta)$. Moreover, $\alpha_{d}<\beta_{d}$ ($\ast$), otherwise $\alpha\prec\beta$ cannot be true. Hence,
$\kappa_{d}(\alpha)$ $<\kappa_{d}(\beta)$
$\overset{\kappa}{\iff}$ $\alpha_{d}+(L_{d}+1)\kappa_{d-1}(\alpha)$ $<\beta_{d}+(L_{d}+1)\kappa_{d-1}(\beta)$
[…]
which is true due to ($\ast$).
Case
$\langle\alpha_{i}\rangle_{i\in[d-1]}\prec\langle\beta_{i}\rangle_{i\in[d-1]}$. In this case, there is no relation between $\alpha_{d}$ and $\beta_{d}$. [^8] To see why this is the case, note that in an English dictionary “root” $\prec$ “rope”, even though the last letter of “rope”, “e”, comes before the “t” of “root”. Hence, we can apply IH2 and use the fact that the difference $\kappa_{d-1}(\beta)-\kappa(\alpha)$ is strictly greater than 0:
$\kappa_{d}(\alpha)$ $<\kappa_{d}(\beta)$
$\overset{\kappa}{\iff}$ $\alpha_{d}+(L_{d}+1)\kappa_{d-1}(\alpha)$ $<\beta_{d}+(L_{d}+1)\kappa_{d-1}(\beta)$
$\iff$ $\alpha_{d}-\beta_{d}$ $<(L_{d}+1)\left(\kappa_{d-1}(\beta)-\kappa_{d-1}(\alpha)\right)$
$\overset{\text{IH2}}{\iff}$ $\alpha_{d}-\beta_{d}$ $<L_{d}+1,$
which is true since the difference $\alpha_{d}-\beta_{d}$ can only lie in $L_{d},L_{d}]$.
This completes the injectivity argument, hence $\kappa$ is a monomorphism. ∎
Appendix B Edge-Weighted $k$-Clique with Small Weights
We show that Nešetřil-Poljak’s algorithm [uw-k-clique] can be extended to support small weights. We consider the $3k$-case here; both the $3k+1$- and $3k+2$-cases are handled analogously.
{restatable}
[]lemmaNP The Edge-Weighted $3k$-Clique problem on undirected graphs $G=(V,E)$ with integer edge weights in $[0,W]$ can be solved in time $O^{*}(n^{\omega k}\,W)$.
Proof.
We extend Nešetřil–Poljak’s construction to the edge-weighted setting. Let $H$ be the auxiliary graph whose vertices are all $k$-cliques $Y\subseteq V$ of $G$, and where distinct $Y_{1},Y_{2}$ are adjacent iff $Y_{1}\cup Y_{2}$ induces a $2k$-clique in $G$. Hence, $H$ has $N=\Theta(n^{k})$ vertices.
Denote by $w(e)\in[0,W]$ the weight of an edge $e\in E(G)$. For a $k$-clique $Y$, set $w^{\prime}(Y)=w(E(G[Y]))$, the “internal” weight of node $Y$, and for adjacent $Y_{1},Y_{2}$ in $H$, set
$w^{\prime}(Y_{1},Y_{2})=\sum_{\begin{subarray}{c}u\in Y_{1}\\ v\in Y_{2}\end{subarray}}w(uv),$
representing the “cross” weight between $Y_{1}$ and $Y_{2}$. If $A,B,C$ form a triangle in $H$, the total weight of the corresponding $3k$-clique $A\cup B\cup C$ is
$W_{\mathrm{clique}}=w^{\prime}(A)+w^{\prime}(B)+w^{\prime}(C)+w^{\prime}(A,B)+w^{\prime}(B,C)+w^{\prime}(C,A).$
Now, since we would like to have edge weights only (in order to not have a specialized algorithm for coping with both cases), redefine the weight of the edge $\{Y_{1},Y_{2}\}$ as:
$w^{\prime}(Y_{1},Y_{2})=2\sum_{\begin{subarray}{c}u\in Y_{1}\\ v\in Y_{2}\end{subarray}}w(uv)+w^{\prime}(Y_{1})+w^{\prime}(Y_{2}),$
that is, we double the previously edge weight, and add the internal weights of the two nodes. Hence, for every triangle $\{A,B,C\}$ in $H$, we have
$w^{\prime}(A,B)+w^{\prime}(B,C)+w^{\prime}(C,A)=2W_{\mathrm{clique}},$
so minimizing triangle weight in $H$ is equivalent to minimizing $3k$-clique weight in $G$.
Next, note that the minimizing edge weight on $H$ satisfies
$W_{H}^{\max}\leq\big(2k^{2}+2\textstyle{\binom{k}{2}}\big)\,W=(3k^{2}-k)W=\Theta(k^{2}W).$
A minimum-weight triangle in an $N$-vertex edge-weighted graph with integer weights in $0,W_{H}^{\max}]$ can be found via a single distance $(\min,+)$-product in time $\tilde{O}(W_{H}^{\max}\,N^{\omega})$. Instantiating $N=\Theta(n^{k})$ and $W_{H}^{\max}=\Theta(k^{2}W)$ yields time $O^{*}(n^{\omega k}\,W)$. Finally, explicitly constructing $H$ and all $w^{\prime}(\cdot,\cdot)$ values takes $O(n^{2k}\operatorname{poly}(k)\log W)$-time, which is dominated by $n^{\omega k}$ since $\omega>2$. ∎

### v2

@@
-Mind the Gap. Doubling Constant Parametrization of Weighted Problems: TSP, Max-Cut, and More
+Mihail Stoian
 Abstract
@@
 Indeed, there are many problems that resist improvements over the standard algorithm.
-Yet, we have arguments to be optimistic: The amount of work in this direction has led to impressive speedups for well-studied problems such as Max-Cut [maxcut-williams], Hamiltonicity [hamiltonicity, below_all_subsets1], Sched [below_all_subsets2], Independent Set [below_all_subsets3], Dominating Set [below_all_subsets3, below_all_subsets4], and Min $k$-Cut [min_k_cut].
+Yet, we have arguments to be optimistic: The amount of work in this direction has led to impressive speedups for well-studied problems such as Max-Cut [Wil05], Hamiltonicity [Bjö10, Bjö16], Sched [CPPW14], Independent Set [FGK09], Dominating Set [FGK09, vRNvD09], and Min $k$-Cut [LSS23].
 To some extent, these speedups have been achieved mainly for what we will refer to as unweighted problems.
@@
 A concrete example is that of the Traveling Salesman Problem (TSP), which asks to find the shortest tour through $n$ cities.
-The textbook algorithm in $O(2^{n}n^{2})$-time remains, up to some modest polynomial improvements [chan_shaving_logs], the best up to date for the general setting [bellman_dp, held_karp_dp, nederlof_tsp].
-In contrast, its unweighted counterpart, the Hamiltonicity problem, has been (non-trivially) sped up to $O^{*}(1.66^{n})$-time by Björklund [hamiltonicity].
+The textbook algorithm in $O(2^{n}n^{2})$-time remains, up to some modest polynomial improvements [Cha13], the best up to date for the general setting [Bel62, HK61, Ned20].
+In contrast, its unweighted counterpart, the Hamiltonicity problem, has been (non-trivially) sped up to $O^{*}(1.66^{n})$-time by Björklund [Bjö10].
 [^2] The $O^{*}$-notation hides polynomial factors in the input size.
 Interestingly, virtually every new result on the unweighted version of a weighted problem has its own theorem stating that the original problem can be solved in the time of its unweighted counterpart, yet with an extra multiplicative pseudo-polynomial dependence on the largest input weight; this is known as the “(polynomial) embedding technique”.
-Notably, this situation also concerns TSP: Björklund proved that Hamiltonicity can be leveraged to solve TSP by paying the price of a near-linear time-factor depending on the largest edge weight $W$ [hamiltonicity, Thm.
+Notably, this situation also concerns TSP: Björklund proved that Hamiltonicity can be leveraged to solve TSP by paying the price of a near-linear time-factor depending on the largest edge weight $W$ [Bjö10, Thm.
 3].
 This becomes impractical for arbitrarily large $W$.
-The same holds true for the weighted Max-Cut problem: its unweighted counterpart can be solved in $O^{*}(2^{\omega n/3})$-time [maxcut-williams, koivisto-correct], where $\omega<2.371339$ is the matrix multiplication exponent [curr-omega], while its weighted version has to “accept” the additional pseudo-polynomial factor.
+The same holds true for the weighted Max-Cut problem: its unweighted counterpart can be solved in $O^{*}(2^{\omega n/3})$-time [Wil05, Koi06], where $\omega<2.371339$ is the matrix multiplication exponent [ADV+25], while its weighted version has to “accept” the additional pseudo-polynomial factor.
 On the other hand, if the weights are polynomially bounded, then the problematic factor becomes polynomial, and the running time of the weighted version matches, in the $O^{*}$-sense, that of the unweighted version.
@@
 Given the ubiquity of the embedding technique, it is natural to ask whether the gap between the running times of the weighted and unweighted versions of a problem is overcome only in the regime of polynomially bounded weights.
-A negative answer would be consistent with the widespread adoption of the embedding approach, with its inherent pseudo-polynomial dependence [maxcut-williams, koivisto-correct, fsc, set_partitioning, hamiltonicity, saving_space, Cygan_book].
+A negative answer would be consistent with the widespread adoption of the embedding approach, with its inherent pseudo-polynomial dependence [Wil05, Koi06, BHKK07, BHK09, Bjö10, LN10, CFK+15].
 In contrast, a positive answer would re-open the door to a more systematic re-examination of the weighted vs.
-unweighted gap, in the spirit of the “Losing Weights by Gaining Edges” technique for node-weighted problems [losing-ws] or in terms of approximation guarantees [apx-ws].
+unweighted gap, in the spirit of the “Losing Weights by Gaining Edges” technique for node-weighted problems [ALW14] or in terms of approximation guarantees [CST01].
 1.1 Our Results
@@
 The set of input weights $A$ is said to have small doubling if $|A+A|\leq\mathcal{C}|A|$ for some constant $\mathcal{C}>1$, referred to as the doubling constant.
-To maintain notation consistency with prior work on unweighted problems [esa-paper], we prefix problem names with $\mathcal{C}$-” if the set of input weights has doubling constant $\mathcal{C}$.
+To maintain notation consistency with prior work on unweighted problems [RW24], we prefix problem names with $\mathcal{C}$-” if the set of input weights has doubling constant $\mathcal{C}$.
 We also write $O_{\mathcal{C}}$ to note that we suppressed factors that depend only on $\mathcal{C}$.
@@
 {restatable}
-[Meta-algorithm]theoremMETATHM If problem $\mathbf{P}_{w}$ satisfies property $\phi$ and the unweighted version can be solved in time $O(T(n))$, then $\mathcal{C}$-$\mathbf{P}_{w}$ can be solved in time $O_{\mathcal{C}}^{*}(T(n))$.
+[Meta-algorithm]theoremMETATHM If problem $\mathbf{P}_{w}$ satisfies property $\phi$ and the unweighted version can be solved in time $O(T(n))$ by an algebraic algorithm $\mathcal{A}$, then $\mathcal{C}$-$\mathbf{P}_{w}$ can be solved in time $O_{\mathcal{C}}^{*}(T(n))$.
 The following corollaries follow:
@@
 The missing piece is a decomposition of the weights into constant-size coefficient-tuples coming from the generalized arithmetic progression $A$ is a subset of (cf.
-Freiman’s theorem [freiman1]).
+Freiman’s theorem [Fre73]).
 To be able to apply this to any bounded-input algorithm out of the box, we must ensure that the coefficient representation preserves the total order of the original weights.
@@
 Additive Combinatorics in Algorithm Design.
-Our work continues the effort to consider important results from additive combinatorics [tao_vu] as a new toolbox for algorithm design, a connection highlighted in works such as Ref.
-[trevisan_ac, ac_cs_view3, ac_cs_view1, ac_cs_view2].
-One of the first seminal results is that of Chan and Lewenstein [chan-lewenstein], who used the (constructive) Balog-Szemerédi-Gowers (BSG) theorem to speed up the monotone variant of the min-plus convolution on small integers.
-Recently, this has been further used and extended in #APSP [fredman_meets_domi], 3SUM [3sum_sidon], listing 4-cycles [abf], and approximate distance oracles [abf].
+Our work continues the effort to consider important results from additive combinatorics [TV06] as a new toolbox for algorithm design, a connection highlighted in works such as Ref.
+[Tre09, Vio11, Bib13, Lov17].
+One of the first seminal results is that of Chan and Lewenstein [CL15], who used the (constructive) Balog-Szemerédi-Gowers (BSG) theorem to speed up the monotone variant of the min-plus convolution on small integers.
+Recently, this has been further used and extended in #APSP [CVX23], 3SUM [JX23], listing 4-cycles [ABF23], and approximate distance oracles [ABF23].
 We observe that the constructive version of a fundamental theorem in additive combinatorics bears fruit in algorithm design.
-To this end, Randolph and Węgrzycki [esa-paper] made Freiman’s theorem [freiman1], another fundamental result in additive combinatorics, constructive.
+To this end, Randolph and Węgrzycki [RW24] made Freiman’s theorem [Fre73], another fundamental result in additive combinatorics, constructive.
 Their main applications were to SubsetSum and ILP-Feasibility, which have an inherent additive structure.
@@
 Repurposing the unweighted algorithm for the weighted version of a problem is a common theme in work that improves over the textbook solution, as in seminal results such as Ref.
-[maxcut-williams, fsc, set_partitioning, hamiltonicity]; notably, in all these examples, the weighted problem operates over the $(\min,+)$ and $(\max,+)$ semirings.
+[Wil05, BHKK07, BHK09, Bjö10]; notably, in all these examples, the weighted problem operates over the $(\min,+)$ and $(\max,+)$ semirings.
 The key idea is to represent the input weights of the problem instance as monomials so that the bounded-input algorithm can work in the $(+,\times)$ ring.
@@
 In fact, there is a simple exercise to turn any bounded-input algorithm for a problem into an exponential-time $(1+\varepsilon)$-approximation algorithm.
-For the Max-Cut problem, Williams mentions this idea in passing [maxcut-williams], inspired by the literature on the all-pairs shortest paths problem [zwick_apsp]; see, e.g., Ref.
-[approx-min-sum-conv] for more such examples.
+For the Max-Cut problem, Williams mentions this idea in passing [Wil05], inspired by the literature on the all-pairs shortest paths problem [Zwi02]; see, e.g., Ref.
+[Sto24] for more such examples.
 It is interesting to ask whether the doubling constant could play a role in approximation algorithms.
@@
 Furthermore, we use the $O_{\mathcal{C}}$-notation, already established in Ref.
-[esa-paper], to denote that we suppressed factors depending on $\mathcal{C}$, the doubling constant on which we parameterize the time-complexities.
+[RW24], to denote that we suppressed factors depending on $\mathcal{C}$, the doubling constant on which we parameterize the time-complexities.
 For instance, $O_{\mathcal{C}}(2^{n})=f(\mathcal{C})\cdot O(2^{n})$ for a computable function $f$.
@@
 For a treatment on the definition of a combinatorial algorithm, see Ref.
-[fast-bmm].
+[AFK+24].
 Due to the additive structure of the input weights we will be exploiting, starting in Sec.
@@
 Freiman’s Theorem.
-Given a set $A$ of small doubling, Freiman’s theorem [freiman1] tells that, indeed, $A$ is included in a GAP $G$, whose dimension is independent on $|A|$.
-In this work, we will need the constructive Freiman’s theorem, recently designed by Randolph and Węgrzycki [esa-paper]:
-Theorem 2.2 (Constructive Freiman’s Theorem [esa-paper]).
+Given a set $A$ of small doubling, Freiman’s theorem [Fre73] tells that, indeed, $A$ is included in a GAP $G$, whose dimension is independent on $|A|$.
+In this work, we will need the constructive Freiman’s theorem, recently designed by Randolph and Węgrzycki [RW24]:
+Theorem 2.2 (Constructive Freiman’s Theorem [RW24]).
 Let $A$ be a set of $n$ integers with $|A+A|\leq\mathcal{C}|A|$.
@@
 Specifically, the theorem provides the values $x_{1},x_{2},\dots,x_{d(\mathcal{C})}$ and $L_{1},L_{2},\dots L_{d(\mathcal{C})}$.
-The above theorem gives us a GAP of dimension $d(\mathcal{C})=2^{\mathcal{C}^{O(1)}}$ and volume $v(\mathcal{C})=2^{2^{\mathcal{C}^{O(1)}}}n$, as in the original Freiman’s theorem [freiman1].
-Noteworthy, later works do optimize $d$ and $v$ [freiman-better-1, freiman-better-2, freiman-better-3, freiman-better-4], yet these improvements are not captured by the constructive version.
+The above theorem gives us a GAP of dimension $d(\mathcal{C})=2^{\mathcal{C}^{O(1)}}$ and volume $v(\mathcal{C})=2^{2^{\mathcal{C}^{O(1)}}}n$, as in the original Freiman’s theorem [Fre73].
+Noteworthy, later works do optimize $d$ and $v$ [Cha02, Sch11, San12a, San12b], yet these improvements are not captured by the constructive version.
 This completes the necessary preliminaries on additive combinatorics.
@@
 This enables running times in the regime of the unweighted version of the problem for a large suite of problems, such as TSP, Weighted Max-Cut, Edge-Weighted $k$-Clique, and Minimum Steiner Tree.
-We stress that, compared to prior work [esa-paper], we consider weighted problems that operate over tropical semirings.
+We stress that, compared to prior work [RW24], we consider weighted problems that operate over tropical semirings.
 4.1 TSP Meets Small Doubling
@@
 Condition 1 holds since a tour’s weight is the sum of the weights of its constituent edges.
-Next, we can take $\mathcal{A}$ as Björklund’s algorithm [hamiltonicity] for undirected Hamiltonicity, which runs in time $O^{*}(1.66^{n})$, and solves the weighted case in time $O^{*}(1.66^{n}\,W)$ [hamiltonicity, Thm.
+Next, we can take $\mathcal{A}$ as Björklund’s algorithm [Bjö10] for undirected Hamiltonicity, which runs in time $O^{*}(1.66^{n})$, and solves the weighted case in time $O^{*}(1.66^{n}\,W)$ [Bjö10, Thm.
 3].
@@
 Another well-studied problem whose unweighted version has been sped up and fits the small doubling setting well is the Weighted Max-Cut problem.
-In a breakthrough result, Williams [maxcut-williams] improved the time complexity of the unweighted Max-Cut to $O^{*}(2^{\omega n/3})$.
+In a breakthrough result, Williams [Wil05] improved the time complexity of the unweighted Max-Cut to $O^{*}(2^{\omega n/3})$.
 We show that we can leverage this result to the weighted setting when the edge weights form an integer set that has small doubling.
@@
 Condition 1 holds since a cut’s weight is the sum of the weights of its constituent edges.
-Next, we can take $\mathcal{A}$ as Williams’s algebraic algorithm [maxcut-williams] (and later improved by Koivisto [koivisto-correct]) for the unweighted Max-Cut problem, which runs in time $O^{*}(2^{\omega n/3})$, and solves the weighted case in time $O^{*}(2^{\omega n/3}\,W)$ [koivisto-correct, Thm.
+Next, we can take $\mathcal{A}$ as Williams’s algebraic algorithm [Wil05] (and later improved by Koivisto [Koi06]) for the unweighted Max-Cut problem, which runs in time $O^{*}(2^{\omega n/3})$, and solves the weighted case in time $O^{*}(2^{\omega n/3}\,W)$ [Koi06, Thm.
 3].
@@
 4.3 Small Doubling and Edge-Weighted $k$-Cliques
-We now turn to a problem whose complexity is a rather popular hypothesis in hardness proofs [clique-hyp-1, clique-hyp-2, clique-hyp-3]: the edge-weighted $k$-clique problem.
-Compared to the problems considered so far, the algebraic algorithm for the unweighted $k$-Clique has already been known from the 80’s [uw-k-clique].
+We now turn to a problem whose complexity is a rather popular hypothesis in hardness proofs [BT17, AVW14, LVW18]: the edge-weighted $k$-clique problem.
+Compared to the problems considered so far, the algebraic algorithm for the unweighted $k$-Clique has already been known from the 80’s [NP85].
 We show that it can be leveraged in the doubling constant parametrization of the edge-weighted version.
 Note that, apart from a mention in passing in Lincoln et al.
-[clique-hyp-3] that Nešetřil-Poljak’s algorithm can also be used for polynomially bounded weights, we could not find any reference to a proof.
+[LVW18] that Nešetřil-Poljak’s algorithm can also be used for polynomially bounded weights, we could not find any reference to a proof.
 For completeness, we show this in Lemma B (Appendix B).
@@
 Condition 1 holds since a clique’s weight is the sum of the weights of its constituent edges.
-Next, we can take $\mathcal{A}$ as Nešetřil-Poljak’s algebraic algorithm [uw-k-clique] for the unweighted version of the problem (or the improvement by Eisenbrand and Grandoni [faster-uw-k-clique] for some values of $k$ using rectangular matrix multiplication), which runs in time $O^{*}(n^{k\omega/3})$, and solves the weighted case in time $O^{*}(n^{k\omega/3}\,W)$ (Lemma B).
+Next, we can take $\mathcal{A}$ as Nešetřil-Poljak’s algebraic algorithm [NP85] for the unweighted version of the problem (or the improvement by Eisenbrand and Grandoni [EG04] for some values of $k$ using rectangular matrix multiplication), which runs in time $O^{*}(n^{k\omega/3})$, and solves the weighted case in time $O^{*}(n^{k\omega/3}\,W)$ (Lemma B).
 Notably, due to its design, namely that each original input weight is doubled in the auxiliary graph, the bounded-input algorithm (Lemma B) has a $\lambda\leq 2|w(I)|$.
@@
 4.4 Steiner Meets Freiman
-For reference, we also include the Minimum Steiner Tree, even though the general setting can be directly solved in $O((2+\epsilon)^{k}\,p(n))$ [faster_steiner1], where $p(n)$ is a polynomial function of $n$, the degree of which grows rapidly as $\epsilon$ approaches zero, Its degree has been later refined to $12\sqrt{\epsilon^{-1}\ln\epsilon^{-1}}$, having as a result running times such as $O(2.1^{k}\,n^{57.6})$ or $O(2.5^{k}\,n^{14.2})$ [fuchs2007dynamic].
-Notably, Dreyfus and Wagner [dreyfus1971steiner] were the first to design a dynamic programming recursion running in time $O(3^{k}n+2^{k}n^{2}+nm)$, where $n=|V|$, $m=|E|$, and $k=|T|$.
+For reference, we also include the Minimum Steiner Tree, even though the general setting can be directly solved in $O((2+\epsilon)^{k}\,p(n))$ [MRR06], where $p(n)$ is a polynomial function of $n$, the degree of which grows rapidly as $\epsilon$ approaches zero, Its degree has been later refined to $12\sqrt{\epsilon^{-1}\ln\epsilon^{-1}}$, having as a result running times such as $O(2.1^{k}\,n^{57.6})$ or $O(2.5^{k}\,n^{14.2})$ [FKM+07].
+Notably, Dreyfus and Wagner [DW71] were the first to design a dynamic programming recursion running in time $O(3^{k}n+2^{k}n^{2}+nm)$, where $n=|V|$, $m=|E|$, and $k=|T|$.
 The doubling constant parametrization of the problem is the following:
@@
 Condition 1 holds since, as the edge weights are all non-negative, the optimal subgraph $H$ is a tree with its leaves in $K$, and a tree’s weight is the sum of the weights of its constituent edges.
-Next, we can take $\mathcal{A}$ as the fast subset convolution based algorithm due to Björklund, Husfeldt, Kaski, and Koivisto [fsc, Sec.
-4.1.2], which runs in $O^{*}(2^{k}n^{2}W)$-time, or Lokshtanov–Nederlof’s polynomial-space algorithm [saving_space] in the same running time.
+Next, we can take $\mathcal{A}$ as the fast subset convolution based algorithm due to Björklund, Husfeldt, Kaski, and Koivisto [BHKK07, Sec.
+4.1.2], which runs in $O^{*}(2^{k}n^{2}W)$-time, or Lokshtanov–Nederlof’s polynomial-space algorithm [LN10] in the same running time.
 ∎
@@
 A simple application is that of the $(\min,+)$-convolution: Given sequences $(a[i])_{i=0}^{n-1}$ and $(b[i])_{i=0}^{n-1}$, the goal is to compute $(c[i])_{i=0}^{n-1}$, where $c[k]=\min_{i=0,\ldots,k}\{a[i]\,+\,b[k-i]\}$.
-The naive algorithm runs in $O(n^{2})$-time, and no $O(n^{2-\varepsilon})$-time algorithm is known for $\varepsilon>0$ [min-plus-1, min-plus-2].
+The naive algorithm runs in $O(n^{2})$-time, and no $O(n^{2-\varepsilon})$-time algorithm is known for $\varepsilon>0$ [CMWW17, KPS17].
 However, the convolution in the $(+,\times)$-ring can be solved in $O(n\log n)$-time via FFT.
 This resembles the setting we considered throughout the paper.
-Indeed, Węgrzycki outlines in his PhD thesis a bounded-input algorithm running in time $\widetilde{O}(nW)$, where $W$ is the largest input value [karol_phd, Lemma 5.7.2].
+Indeed, Węgrzycki outlines in his PhD thesis a bounded-input algorithm running in time $\widetilde{O}(nW)$, where $W$ is the largest input value [Węg19, Lemma 5.7.2].
 Applying the same key idea as for the NP-hard problems we considered, we can obtain an $\widetilde{O}_{\mathcal{C}}(n)$-time algorithm for the min-plus self-convolution problem, i.e., $a=b$, and the input sequence, when regarded as a set, has small doubling.
-The same setting holds true for the min-sum subset convolution [fsc].
-It is thus interesting to ask how the framework behaves in more weighted applications, such as (specialized) $(\min,+)$ matrix products [special-min-plus-prod] or even APSP, where—similar to the NP-hard case studied here—the objective value is an additive combination of input weights.
+The same setting holds true for the min-sum subset convolution [BHKK07].
+It is thus interesting to ask how the framework behaves in more weighted applications, such as (specialized) $(\min,+)$ matrix products [VX20] or even APSP, where—similar to the NP-hard case studied here—the objective value is an additive combination of input weights.
 Acknowledgments.
@@
 Appendix B Edge-Weighted $k$-Clique with Small Weights
-We show that Nešetřil-Poljak’s algorithm [uw-k-clique] can be extended to support small weights. We consider the $3k$-case here; both the $3k+1$- and $3k+2$-cases are handled analogously.
+We show that Nešetřil-Poljak’s algorithm [NP85] can be extended to support small weights. We consider the $3k$-case here; both the $3k+1$- and $3k+2$-cases are handled analogously.
 {restatable}